OpenStudy (anonymous):

sqrt x + 33 plus sqrt x - 2 =7

OpenStudy (anonymous):

I take it you need to solve for x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Organize your work with parentheses first.

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

Erm, its hard to see what your trying to say, because is it \[\sqrt{x} - 2 \] or \[\sqrt{x-2}\] They're two different values, by the way.

OpenStudy (anonymous):

Considering you mean \[\sqrt{x+33}+\sqrt{x-2}=7\] You multiply by \[\sqrt{x+33}-\sqrt{x-2}\] And you get \[(x+33)-(x-2)=35=7(\sqrt{x+33}-\sqrt{x-2})\leftarrow \rightarrow \sqrt{x+33}-\sqrt{x-2}=5\] So you have a system \[\sqrt{x+33}+\sqrt{x-2}=7 \] \[\sqrt{x+33}-\sqrt{x-2}=5\] You add them up and get \[2\sqrt{x+33}=12\] \[\sqrt{x+33}=6\] Meaning \[x+33=36\leftarrow \rightarrow x=3\]

OpenStudy (anonymous):

Let's say its (sqrt x + 33) + (sqrt x - 2) =7

OpenStudy (anonymous):

Oh, u did already...

OpenStudy (anonymous):

Yep I did hehe

OpenStudy (anonymous):

Wolfram agrees with u http://www.wolframalpha.com/input/?i=sqrt+%28x+%2B+33%29+%2B+sqrt+%28x+-+2%29+%3D7